A wheel has a radius of 3.2 m. How far (path length) does a point on the circumference travel if the wheel is rotated through angles of 30°, 30 rad, and 30 rev, respectively?
Answer #1

arc length , l = r \theta

(a) l = 3.2 * pi/6 = 1.675 m

(b) l = 3.2 * 30 = 96 m

(c) l = 3.2 * 30 rev = 3.2 * 30 * 2 * pi = 603.18 m